OpenStudy (anonymous):

Charlie's piggy bank was filled with dimes and nickels. He took the money out and counted a total of 419 coins. The coins were worth $36.05. How many nickels were in the piggy bank?

OpenStudy (anonymous):

x + y = 419 0.05x + 0.1y = 36.05 Solve where x is the number of nickels and y is the number of dimes.

OpenStudy (anonymous):

That's too hard. I really don't know how to solve that.

OpenStudy (anonymous):

How about this? x + y = 419 0.05x + 0.1y = 36.05 x + y = 419 20(0.05x + 0.1y = 36.05) x + y = 419 x + 2y = 721 Can you solve now?

OpenStudy (anonymous):

Yup, there were 117 nickels :)

OpenStudy (anonymous):

Thanks! (:

OpenStudy (anonymous):

np :)